OpenStudy (kanwal32):

The quadratic equation whose root are in A.M and H.M is equal to the sum of squares of their reciprocals, 2x^2-3x+5=0 is (A) 4x^2-3x+5=0 (B)12x^2-49x+30=0 (C)14x^2-12x+35=0 (D)2x^2+3x+5= 0

OpenStudy (kanwal32):

@dan815 pls help

OpenStudy (kanwal32):

@luigi0210 pls save my life

OpenStudy (kanwal32):

@hartnn pls help

hartnn (hartnn):

i am not clear with the question...whose roots are in AM?

OpenStudy (kanwal32):

i will rewrite the question

OpenStudy (kanwal32):

The quadratic equation whose roots are the AM and HM between the roots of the equation 2x^2-3x+5=0

OpenStudy (kanwal32):

is- and option are the same

hartnn (hartnn):

better.

OpenStudy (kanwal32):

i am sorry

OpenStudy (kanwal32):

a bit tensed

hartnn (hartnn):

no problem. :) so, 2x^2-3x+5=0 can you find sum of roots and product of roots of this quadratic ?

OpenStudy (kanwal32):

yes

hartnn (hartnn):

AM means arithmetic mean. AM of 2 numbers will be sum of those 2 numbers divided by 2 so what will be the AM ?

OpenStudy (kanwal32):

wait for a minute

OpenStudy (kanwal32):

but there are no real roots for this equation

hartnn (hartnn):

you can still find the sum of roots. have you ever head that sum if roots for \(ax^2+bx+c\) would be \(\Large \dfrac{-b}{a}\) ?

OpenStudy (kanwal32):

no

OpenStudy (kanwal32):

ok

OpenStudy (kanwal32):

so 3/2

hartnn (hartnn):

correct.

hartnn (hartnn):

now product of roots, for ax^2 +bx+c, product of roots is just c/a

OpenStudy (kanwal32):

ok

OpenStudy (kanwal32):

yes i have heard

OpenStudy (kanwal32):

5/4

OpenStudy (kanwal32):

alpha+beta=-b/a alpha+beta=c/a

hartnn (hartnn):

thats correct! so, AM = sum/2 = 3/4

hartnn (hartnn):

know what HM is ?

OpenStudy (kanwal32):

sqrtab

hartnn (hartnn):

no, thats geometric mean :)

OpenStudy (kanwal32):

ohhh sorry

hartnn (hartnn):

1/HM = 1/a + 1/b

OpenStudy (kanwal32):

yes

hartnn (hartnn):

so HM = ab/ (a+b)

hartnn (hartnn):

ab = product = 5/2 a+b = sum = 3/2 find HM now

OpenStudy (kanwal32):

(a+b/n)^-1

OpenStudy (kanwal32):

5/3

hartnn (hartnn):

yep, HM = 5/3 so now you have both roots 3/4 and 5/3 can you make an equation ?

OpenStudy (kanwal32):

yes

OpenStudy (kanwal32):

so sorry i asked such a silly question

hartnn (hartnn):

i wasn't silly. it was a good question :) and even silly questions are welcome here, there's always something to learn from silly questions too :)

hartnn (hartnn):

**it wasn't silly

OpenStudy (kanwal32):

thnx

OpenStudy (kanwal32):

i read it gm

hartnn (hartnn):

welcome ^_^